OpenStudy (anonymous):

Find an equation of the line satisfying the given conditions passing through (2,-9) and perpendicular to the line 5y = -(2x/9) + 5

OpenStudy (anonymous):

Find the slope m of 5y = -2x/9 + 5 by dividing both side for 5 Can you do it?

OpenStudy (anonymous):

how will that get the equation?

OpenStudy (anonymous):

first:get the derivative of your equation

OpenStudy (anonymous):

just do whatever Chlorophyll told you.. :D..

OpenStudy (anonymous):

can someone please give me the answer i honestly have no clue

OpenStudy (anonymous):

1

OpenStudy (anonymous):

-10/9

OpenStudy (anonymous):

(-2x+45)/45

OpenStudy (anonymous):

Divide 5 to the right side -(2x/9) -> -(2x/45) + 1

OpenStudy (anonymous):

Do you know slope m now?

OpenStudy (anonymous):

m=1 ?

OpenStudy (anonymous):

y = mx + b Try again!

OpenStudy (anonymous):

what is that? ^

OpenStudy (anonymous):

maths is my worst subject i am very sorry

OpenStudy (anonymous):

This is linear equation in slope- intercept form (m is slope, b is y intercept) y = mx + b y = -(2x/45) + 1 Now try to find slope m this time

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Matching to find what number in front of x, can you?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

No, the whole faction !

OpenStudy (anonymous):

slope = -2/45 ?

OpenStudy (anonymous):

There you go!

OpenStudy (anonymous):

Now, can you find the slope of perpendicular line?

OpenStudy (anonymous):

45/2 ?

OpenStudy (anonymous):

now?

OpenStudy (anonymous):

No, flip it 2/45

OpenStudy (anonymous):

so is this the final answer? y = -(2x/45) + 1

OpenStudy (anonymous):

NO, you use the given linear line to find slope of the perpendicular line! => m = 2/45

OpenStudy (anonymous):

Now the slope and give point into Point Slope formula: y - y1 = m ( x - x1) Can you?

OpenStudy (anonymous):

i have no clue now

OpenStudy (anonymous):

given point (x1, y1) = (2, -9)